OpenStudy (anonymous):

What is the sum of the first five terms of a geometric series with a1 = 20 and r = 1/4?

OpenStudy (anonymous):

G.S : a/(1-r) =20/(1-0.25) =26.67

OpenStudy (campbell_st):

thats the limiting sum... you need \[S_{n} = \frac{a(1- r^n)}{1 - r}\] a = 20, r = 0.25 and n = 5